emoneutral:

please i really need help with this The interior angles formed by the sides of a regular nonagon (a polygon with 9 sides) have measures that sum to 405 degrees. What is the degree measurement of each angle? (a) Draw a nonagon. Number each side, (b) Let n represent the measure of an angle in this regular nonagon. What is the degree measure of each angle? Show your work

surjithayer:

sum of interior angles of a nonagon=(9-2)*180=7*180=1260 degrees so there is something wrong in your statement.

surjithayer:

sum of exterior angles of a regular polygon=360 each exterior angle=360/9=40 degree so each interior angles of a regular nonagon=180-40=140 degrees.